# CMOB: Large-Scale Cancer Multi-Omics Benchmark with Open Datasets, Tasks, and Baselines

### v1

@@ A.1 CMOB Structure
│   ├── Pan-Cancer Dataset [Labeled]
@@
│   ├── Cancer Subtype Datasets [Unlabeled]
@@
│   ├── Golden-Standard Cancer Subtype Datasets [Labeled]
@@
│   ├── Omics Data Imputation Datasets [Unlabeled]
@@
│   ├── Classification and Clustering Tasks
@@
└── Complementary Resources
    ├── Mapping File
    │   └── CMOB_STRING_mapping.csv
@@
    └── Biological Attribute Analysis Tools
        └── Analysis_Tools.py
@@ A.2 Dataset Format
CMOB uses .csv files to manage and store all omics datasets, widely favored in biomedical research, including multi-omics studies. .csv files are plain-text files where data is separated by commas. They maintain a straightforward structure, with rows representing individual data records and columns representing different attributes or variables. Despite its simplicity, .csv remains efficient even with large datasets encountered in genomic and proteomic studies.
For example, consider a simplified .csv file containing mRNA data for a set of gene features (rows) across several patient samples (columns):
@@
Both Python and R provide built-in functions and libraries to read, write, and analyze .csv files.
@@ A.3 Recruited Cancer
The CMOB database contains multi-omics data for 32 types of cancer. The full names and abbreviations as shown in the following table:
Table 4: Cancer types and abbreviations in CMOB
@@ A.4 Recruited Omics
CMOB recruited four types of omics data:
@@
Appendix B CMOB Pipelines
B.1 Recruited Omics
B.2 Omics Data Processing Pipelines
@@
(i) For transcriptomics (mRNA and miRNA) data:
@@
    STEP 3: Convert Gene-Level Estimates For data from the Hi-Seq platform like Illumina, use the R package edgeR Robinson et al. (2010) to convert the scaled estimates in the original gene-level RSEM to FPKM.
@@
    STEP 4: Filter Non-Human miRNA For "miRNA-Seq" data from Illumina GA and Agilent array platforms, identify and remove non-human miRNA expression features using species annotation from databases like miRBase Kozomara et al. (2019).
@@
    STEP 3: Identify Recurrent Alterations Use the R package GAIA al. SMe (2021) to identify recurrent alterations in the cancer genome from raw data that denote all aberrant regions resulting from copy number variation segmentation.
@@
    STEP 4: Annotate Genomic Regions Use the R package BiomaRt Durinck et al. (2005) to annotate the aberrant recurrent genomic regions.
@@
(iii) For epigenomic (Methy) data:
@@
    STEP 2: Normalize Methylation Data Implement a median-centering normalization to account for systematic biases and technical variations across samples using the R package limmaRitchie et al. (2015).
@@
Appendix C CMOB Tasks
C.1 Pan-cancer Classification
@@
C.2 Cancer Subtype Clustering
Cancer subtyping means categorizing patients into subgroups that exhibit differences in various aspects based on their multi-omics data. However, for most cancer types, especially rare cancers, the cancer subtyping tasks are still open questions under discussion. Thus, cancer subtyping tasks are typically clustering tasks without ground true labels. Let $X^{O}=\{x_{1},x_{2},...,x_{m}\}$ represent the multi-omics dataset, where each $x_{i}$ is a vector of features in $O$-th omics for the $i$-th sample. Let $k$ denote the set of possible cancer subtypes. The goal of cancer subtyping using multi-omics data is to assign each sample $x_{i}$ in $X$ into $k$ clusters $C=\{C_{1},C_{2},...,C_{k}\}$, such that each cluster $C_{i}$ represents a distinct cancer subtype based on the information from multiple omics data sources.
C.3 Golden-standard Subtype Classification
@@
C.4 Omics Data Imputation
@@
Appendix D Evaluation Metrics
D.1 Silhouette coefficient (SIL)
@@
D.2 P-value of the log-rank test on survival time (LPS)
@@
D.3 Precision (PREC)
$PREC=\frac{TP}{TP+FP}$
D.4 Normalized mutual information (NMI)
$NMI(C,G)=\frac{I(C,G)}{\sqrt{H(C)\cdot H(G)}}$
D.5 Adjusted rand index (ARI)
D.6 Mean absolute error (MAE)
@@
D.7 Root mean squared error (RMSE)
@@
Table: The evaluation metrics and number of baselines used for all tasks in CMOB. Tasks #Baselines Metrics Pan-cancer Classification 10 PREC, NMI, ARI Cancer Subtype Clustering 10 SIL, LPS Golden-standard Subtype Classification 10 PREC, NMI, ARI Omics Data Imputation 7 MAE, RMSE
Appendix E Complementary Resources
E.1 Differential Gene Expression Analysis.
Differential gene expression analysis has been a cornerstone of transcriptomic studies. In this analysis, we compare gene expression levels between different experimental conditions or sample groups to identify genes that are significantly upregulated or downregulated. Statistical tests such as t-tests or non-parametric tests are commonly used for this purpose. For example, comparing gene expression profiles between cancer patients and healthy controls to identify genes that are dysregulated in cancer. Genes with significant differences in expression levels may be further investigated as potential biomarkers or therapeutic targets. For example, researchers performed differential gene expression analysis on RNA-seq data from Alzheimer’s disease patients and healthy controls. This analysis identified a panel of differentially expressed genes implicated in neuroinflammation and synaptic dysfunction, showing molecular pathways associated with Alzheimer’s disease progression.
@@ Appendix E Complementary Resources
Among the identified DEGs, several genes have been extensively reported as being associated with cancer progression. Notable examples include BRCA1, WNT4, and NOTCH2. BRCA1 is well-known for its involvement in hereditary breast cancer and plays essential roles in cell cycle regulation, DNA damage response, and transcriptional control Miki et al. (1994). Dysregulation of the WNT4 gene, which encodes a protein belonging to the Wnt signaling pathway, has been linked to tumor growth, invasion, and metastasis Peradziryi et al. (2011). Similarly, the NOTCH2 gene, a member of the Notch receptor family, is critical in cell fate determination, development, and tissue homeostasis and has been implicated in tumor initiation, progression, and therapy resistance Gridley (2003).
E.2 Pathway Analysis.
Pathway analysis is a critical step in interpreting the biological significance of DEGs. By mapping DEGs to known biological pathways, researchers can gain insights into the underlying mechanisms and potential functional impacts of gene expression changes. In CMOB, pathway analysis is performed using established databases such as KEGG. These databases provide curated information on metabolic pathways, signaling cascades, and gene ontologies. DEGs are input into pathway analysis tools to conduct the analysis, which then identifies overrepresented pathways among the upregulated and downregulated gene sets.
@@
Appendix F Data Source Ethics and Policies
@@
F.1 Human Subjects Protection and Data Access Policies:
@@
F.2 Data Use Certification Agreement:
@@
F.3 Suggested Informed Consent Language for Prospective Collections:
@@
F.4 Sharing Data from Large-scale Biological Research Projects:
@@
F.5 Considerations for Open Release of Genomic Data from Human Cancer Cell Lines:
@@
Appendix G Limitations & Broader Impact

### v2

@@ A.1 MLOmics Structure
  │   ├── Pan-Cancer Dataset [Classification_datasets]
@@
  │   ├── Cancer Subtype Datasets [Clustering_datasets]
@@
  │   ├── Golden-Standard Cancer Subtype Datasets [Classification_datasets]
@@
  │   ├── Omics Data Imputation Datasets [Imputation_datasets]
@@
  │   ├── Classification Tasks
  │   │   ├── Baselines.py
  │   │   └── Metrics.py
  │   ├── Clustering Tasks
@@
  └── Downstream Analysis Tools and Resources Linking
      ├── Knowledge_bases
      │   └── STRING_mapping / KEGG_mapping.csv
@@
      └── Analysis Tools
          └── Analysis_Tools.py/r
@@ A.2 Dataset Format
MLOmics uses CSV files to manage and store all omics datasets, widely favored in biomedical research, including multi-omics studies. CSV files are plain-text files where data is separated by commas. They maintain a straightforward structure, with rows representing individual data records and columns representing different attributes or variables. Despite its simplicity, CSV remains efficient even with large datasets encountered in genomic and proteomic studies.
For example, consider a simplified CSV file containing mRNA data for a set of gene features (rows) across several patient samples (columns):
@@
Both Python and R provide built-in functions and libraries to read, write, and analyze CSV files efficiently.
@@ A.3 Recruited Cancer
The MLOmics database contains multi-omics data for 32 types of cancer. The full names and abbreviations as shown in the following table:
@@ A.4 Recruited Omics
MLOmics recruited four types of omics data:
@@
Appendix B MLOmics Preprocessing Pipelines
@@
B.1 For Transcriptomics (mRNA and miRNA) Data
@@
    STEP 3: Convert Gene-Level Estimates For data from the Hi-Seq platform like Illumina, use the R package edgeR [17] to convert the scaled estimates in the original gene-level RSEM to FPKM.
@@
    STEP 4: Filter Non-Human miRNA For "miRNA-Seq" data from Illumina GA and Agilent array platforms, identify and remove non-human miRNA expression features using species annotation from databases like miRBase [18].
@@ B.2 For Genomic (CNV) Data
    STEP 3: Identify Recurrent Alterations Use the R package GAIA [19] to identify recurrent alterations in the cancer genome from raw data that denote all aberrant regions resulting from copy number variation segmentation.
@@
    STEP 4: Annotate Genomic Regions Use the R package BiomaRt [20] to annotate the aberrant recurrent genomic regions.
@@
B.3 For Epigenomic (Methy) Data
@@
    STEP 2: Normalize Methylation Data Implement a median-centering normalization to account for systematic biases and technical variations across samples using the R package limma[21].
@@
Appendix C MLOmics Feature Scale Processing
Cancer multi-omics analysis often suffers from data issues, such as unbalanced sample sizes and feature dimensions.
To address this, after the typical omics data preprocessing, MLOmics provides three versions of feature scales (Original, Top, and Aligned) to support different machine learning tasks.
C.1 Original Features
The original features are genes directly extracted from each preprocessed omics dataset. It represents the complete, full-size set of patients’ gene features without any task-specific filtering. This version supports users in customizing their datasets based on their specific requirements, such as re-filtering to target gene sets or omics-specific transformations.
Key technical operations in generating original features include:
  1. 1.
    Retaining all genes after normalization (e.g., log transformation or z-score normalization).
  2. 2.
    Imputing missing values using methods like K-nearest neighbors (KNN) or median imputation.
  3. 3.
    Filtering out low-quality samples (e.g., samples of low-variances or high missing genes).
C.2 Aligned Features
Aligned features are the intersection of genes common to all datasets for a learning task, representing the shared features in different cancer types. This reduces the original feature size, ensures consistency in multi-omics features, and primarily provides support across cancer-type studies.
Key technical operations for generating aligned features include:
  1. 1.
    Resolving unmatches in gene naming formats (e.g., ensuring compatibility between cancers using different references genome).
  2. 2.
    Identifying the intersection of feature lists across datasets to ensure all selected features are present in different cancers.
  3. 3.
    Normalization features(e.g., log transformation or z-score normalization).
C.3 Top Features
Top features are selected based on ANOVA [11] statistical testing, ranked by p-values to identify the most significant features across cancers. The default top feature scale settings for mRNA, miRNA, methylation, and CNV data are 5000, 200, 5000, and 5000, respectively, with significance determined by $p<0.05$. This approach greatly reduces noisy genes across cancers and achieves smaller feature-dimension cancer datasets, making them suitable for feature-dimension-sensitive machine-learning models.
Additional key technical operations for generating top features include:
  1. 1.
    Performing multi-class ANOVA to identify genes with significant variance across multiple cancer types.
  2. 2.
    Adjusting for multiple testing using the Benjamini-Hochberg correction to control the false discovery rate (FDR).
  3. 3.
    Ranking features by adjusted p-values and selecting the top $k$ features per omics type, as defined by the default or user-specified scales.
  4. 4.
    Normalization features(e.g., log transformation or z-score normalization).
The detailed feature size of different MLOmics datasets is below:
Table 2: MLOmics provides multiple feature scales for nine unlabeled cancer subtype datasets and five labeled, golden-standard subtype datasets.
Dataset Feature Scale Omics Feature Size
mRNA miRNA Methy CNV
ACC Orignal 18204 368 19045 19525
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
KIRP Orignal 17254 375 19023 19532
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
KIRC Orignal 18464 352 19045 19523
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
LIHC Orignal 17945 435 19053 19523
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
LUAD Orignal 18303 435 19034 19532
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
LUSC Orignal 18577 745 19025 19543
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
PRAD Orignal 17954 467 19034 19534
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
THCA Orignal 17480 345 19024 19532
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
THYM Orignal 18341 535 19034 19532
Aligned 10452 254 10347 10154
Top 5000 200 5000 5000
GS-COAD Orignal 18234 462 19023 19545
Aligned 11343 286 11189 11203
Top 5000 200 5000 5000
GS-BRCA Orignal 18233 345 19053 19533
Aligned 11343 286 11189 11203
Top 5000 200 5000 5000
GS-GBM Orignal 17545 335 19034 19545
Aligned 11343 286 11189 11203
Top 5000 200 5000 5000
GS-LGG Orignal 18345 345 19023 19534
Aligned 11343 286 11189 11203
Top 5000 200 5000 5000
GS-OV Orignal 1735 244 19034 19534
Aligned 11343 286 11189 11203
Top 5000 200 5000 5000
Appendix D MLOmics Tasks
D.1 Pan-cancer Classification
@@
D.2 Cancer Subtype Clustering
Cancer subtyping means categorizing patients into subgroups that exhibit differences in various aspects based on their multi-omics data. However, for most cancer types, especially rare cancers, the cancer subtyping tasks are still open questions under discussion. Thus cancer subtyping tasks are typically clustering tasks without ground true labels. Let $X^{O}=\{x_{1},x_{2},...,x_{m}\}$ represent the multi-omics dataset, where each $x_{i}$ is a vector of features in $O$-th omics for the $i$-th sample. Let $k$ denote the set of possible cancer subtypes. The goal of cancer subtyping using multi-omics data is to assign each sample $x_{i}$ in $X$ into $k$ clusters $C=\{C_{1},C_{2},...,C_{k}\}$, such that each cluster $C_{i}$ represents a distinct cancer subtype based on the information from multiple omics data sources.
D.3 Golden-standard Subtype Classification
@@
D.4 Omics Data Imputation
@@
Appendix E MLOmics Evaluation Metrics
E.1 Precision (Pre)
$Pre=\frac{TP}{TP+FP}$
E.2 Recall (Re)
Recall, also known as sensitivity, measures the ability of a classification or clustering model to identify all relevant instances (i.e., TP) correctly. It is defined as the ratio of TP predictions to the total number of actual positive instances:
$Re=\frac{TP}{TP+FN}$
where $TP$ is the number of true positive predictions (instances correctly classified as positive), and $FN$ is the number of false negative predictions (instances incorrectly classified as negative).
E.3 F1-Score (F1)
The F1-score is the harmonic mean of precision and recall, and it provides a balanced measure of a model’s accuracy by considering both false positives and false negatives. It is particularly useful when the dataset is imbalanced. The F1-score is calculated as:
$F1=2\cdot\frac{Pre\cdot Re}{Pre+Re}$
where $Pre$ is precision, and $Re$ is recall.
E.4 Normalized Mutual Information (NMI)
$NMI(C,G)=\frac{I(C,G)}{\sqrt{H(C)\cdot H(G)}}$
E.5 Adjusted Rand Index (ARI)
E.6 Silhouette Coefficient (SIL)
@@
E.7 P-value of the log-rank Test on Survival Time (LPS)
@@
E.8 Mean Absolute Error (MAE)
@@
E.9 Root Mean Squared Error (RMSE)
@@
Appendix F Downstream Analysis and Biological Resources Linking
F.1 Differential Gene Expression Analysis
Differential gene expression analysis has been a cornerstone of transcriptomic studies. In this analysis, we compare gene expression levels between different experimental conditions or sample groups to identify genes that are significantly upregulated or downregulated. Statistical tests such as t-tests or non-parametric tests are commonly used for this purpose. For example, gene expression profiles between cancer patients and healthy controls can be compared to identify genes that are dysregulated in cancer. Genes with significant differences in expression levels may be further investigated as potential biomarkers or therapeutic targets. For example, researchers performed differential gene expression analysis on RNA-seq data from Alzheimer’s disease patients and healthy controls. This analysis identified a panel of differentially expressed genes implicated in neuroinflammation and synaptic dysfunction, showing molecular pathways associated with Alzheimer’s disease progression.
@@ F.1 Differential Gene Expression Analysis
Among the identified DEGs, several genes have been extensively reported as being associated with cancer progression. Notable examples include BRCA1, WNT4, and NOTCH2. BRCA1 is well-known for its involvement in hereditary breast cancer and plays essential roles in cell cycle regulation, DNA damage response, and transcriptional control [50]. Dysregulation of the WNT4 gene, which encodes a protein belonging to the Wnt signaling pathway, has been linked to tumor growth, invasion, and metastasis [51]. Similarly, the NOTCH2 gene, a member of the Notch receptor family, is critical in cell fate determination, development, and tissue homeostasis and has been implicated in tumor initiation, progression, and therapy resistance [52].
F.2 Survival Analysis
Survival analysis is a vital statistical method used to examine and interpret the time until the occurrence of an event, such as death, disease progression, or relapse, in clinical studies. It provides insights into factors that influence the survival probability of patients and helps in understanding the impact of clinical, demographic, and molecular variables on patient outcomes. Common survival analysis techniques include the Kaplan-Meier estimator for survival curves and the Cox proportional hazards model for assessing the relationship between survival and multiple covariates.
In MLOmics, survival analysis is conducted using time-to-event data, such as patient survival time and event status (alive/dead or disease-free/relapsed). In our approach, we use log-rank tests to compare survival curves between different groups and assess the significance of survival differences. Multivariate Cox regression is employed to evaluate the combined effect of multiple factors on survival. Adjustments for confounding variables and interactions are made, and results are presented with hazard ratios and corresponding confidence intervals. Survival analysis results are visualized using Kaplan-Meier survival curves.
F.3 KEGG Pathway Analysis
Pathway analysis is a critical step in interpreting the biological significance of DEGs. By mapping DEGs to known biological pathways, researchers can gain insights into the underlying mechanisms and potential functional impacts of gene expression changes. In MLOmics, pathway analysis is performed using established databases such as KEGG. These databases provide curated information on metabolic pathways, signaling cascades, and gene ontologies. DEGs are input into pathway analysis tools to conduct the analysis, which then identifies overrepresented pathways among the upregulated and downregulated gene sets.
@@ F.3 KEGG Pathway Analysis
F.4 STRING Network Mapping
The STRING database [12] aggregates PPIs from experimental data, computational predictions, and curated datasets, offering a standardized framework for network analysis. STRING network mapping is used to identify and analyze protein-protein interactions (PPIs) among differentially expressed genes. This approach facilitates the identification of hub nodes and key interaction pathways in different patients and disease groups. For example, patient clusters often correspond to functional modules or biological pathways, leading to different gene networks.
In omics analyses, gene identifiers often differ across databases and platforms, which can pose challenges in integrating data for downstream analyses. Omics datasets may use Ensembl IDs, Entrez IDs, or gene symbols, while the STRING database requires its own set of identifiers to query PPIs. This step is essential for maintaining data consistency and enabling precise network analysis: without proper mapping, some genes might be excluded from the analysis due to identifier mismatches, leading to incomplete or biased results.
In MLOmics, a mapping file resolves these discrepancies by linking MLOmics gene identifiers from omics data to their corresponding STRING identifiers. This mapping file is a CSV format file that contains two columns. The first column is gene identifiers used in the MLOmics dataset. The second column provides the matching STRING identifiers required for querying the STRING database. This structure ensures a straightforward lookup for identifier conversion. Moreover, the CSV format makes inspecting, updating, and adapting this mapping file for other workflows or databases easy.
Once identifiers are mapped, DEGs can be input into the STRING database to construct interaction networks and further network visualization, typically performed with node attributes (e.g., gene expression values or statistical significance) and edge attributes (e.g., interaction confidence) encoded in the visualization.
F.5 Simulate Gene Knockout
The simulation begins by ranking all genes based on node degree disparities calculated from the connectivity matrices of the sub-networks. Node degree is quantified as the number of direct connections each gene has to other genes within the network, serving as a measure of its centrality and influence across different cancer subtypes. To derive the connectivity matrices, we analyze the interactions between genes, where each gene is represented as a node and each interaction as an edge. The degree of each node is then computed to identify highly interconnected genes.
After ranking, we categorize the genes into two sets: a high-ranking gene set, which includes genes exhibiting the largest degree disparities (above a defined threshold based on node degree variance), and a low-ranking gene set, composed of genes with minimal degree differences (below the same threshold). Using node degree variance as a threshold ensures our classification is statistically grounded. This method isolates genes that play critical roles in the network dynamics.
Next, we individually simulate the knockout of genes within the high-ranking and low-ranking gene sets. This process involves transforming their expression values to a baseline non-expression level, which is defined as either zero or a predefined low expression value (such as the mean expression level of the lowest 10% of genes). This transformation mimics the functional loss of these genes. For each gene target in the selected sets, we systematically replace its expression value in the patient samples with the baseline non-expression level.
Appendix G Data Source Ethics and Policies
@@
G.1 Human Subjects Protection and Data Access Policies
@@
G.2 Data Use Certification Agreement
@@
G.3 Suggested Informed Consent Language for Prospective Collections
@@
G.4 Sharing Data from Large-scale Biological Research Projects
@@
G.5 Considerations for Open Release of Genomic Data from Human Cancer Cell Lines
@@
Appendix H Limitations & Broader Impact